a garden has the shape of a right triangle with one leg 5 meters longer than the other. the hypotenuse is 5 meters less than twice the length of the shorter leg
what is the length of the shorter leg?

Answer by mananth(16946)

shorter leg =x
longer leg = x+5
hyp=2x-5
...
Apply Pythagoras theorem
x^2+(x+5)^2=(2x-5)^2
x^2+x^2+10x+25=4x^2-20x+25
group the terms
2x^2-30x=0
2x(x-15)=0
so x= 15
..
15^2+20 ^2= 625
(2x-5)^2= 625